OpenStudy (anonymous):

R,S, and T are collinear, and S is betrween R and T. If RS=x+1, ST=2x-2, and RT=5x-5, find RT.

OpenStudy (anonymous):

 If you look at the (badly drawn) picture, you can see that RS + ST = RT. You can substitute in your equations to solve for x. (x + 1) + (2x - 2) = (5x -5) 3x - 1 = 5x - 5 -2x = -4 x = 2 The question asked to find what RT is. Remember that RT = 5x - 5, plug in x = 5 to that equation. RT = 5(2) - 5 = 10 - 5 = 5 RT = 5